OpenStudy (wilder.monday):

PLEASE HELP 1. 2 log x = 3 log 5

OpenStudy (wilder.monday):

@SolomonZelman

OpenStudy (wilder.monday):

i think the next step would be log x^2 = log 5^3

OpenStudy (wilder.monday):

and then maybe next would be log x^2 = log 125 but i'm not sure about this step or what to do next...

OpenStudy (dannyo19):

log base 10?

OpenStudy (dannyo19):

Or natural log?

OpenStudy (wilder.monday):

log base 10

OpenStudy (wilder.monday):

i think...

OpenStudy (wilder.monday):

@zepdrix

OpenStudy (dannyo19):

So 2logbase10 of x = 3logbase10 of 5

OpenStudy (wilder.monday):

yes

OpenStudy (dannyo19):

Do you know about natural logs?

OpenStudy (wilder.monday):

my teacher tried to teach me but honestly im confused

OpenStudy (dannyo19):

The latter half of the equation is saying 3 times 10 to what exponent equals 5

zepdrix (zepdrix):

Oh sorry I ran off for a bit :D So you need to apply your Log-Exponent Rule: \(\rm b\cdot\log(a)=\log(a^b)\) \[\large\rm 2\log(x)=3\log(5)\]Applying this rule to the left side of the equation gives us,\[\large\rm \log(x^2)=3\log(5)\]Understand how you can apply it to the other side as well? Try it! :)

OpenStudy (dannyo19):

3 times 10^x = 5 == 3logbase10 of 5

OpenStudy (wilder.monday):

log 5^3 @zepdrix

OpenStudy (dannyo19):

So to solve for x divide by 3 than take the natural log of both sides.

zepdrix (zepdrix):

\[\large\rm \log(x^2)=\log(5^3)\]Good good good. Now notice that the logs are the same, so the insides must equal as well,\[\large\rm x^2=5^3\]

OpenStudy (wilder.monday):

so it has to be x^2 = 125

zepdrix (zepdrix):

Yes, then solve for x.

OpenStudy (dannyo19):

or that a way

OpenStudy (wilder.monday):

its a decimal. is that okay?

zepdrix (zepdrix):

I dunno, you'll have to read your instructions for that, exact form, or decimal approximation.

zepdrix (zepdrix):

One thing to note: Usually when you take a square root, you get a plus/minus for solutions.\[\large\rm x=\pm\sqrt{125}\]But notice that x=-sqrt(125) does NOT work for our original problem. You can't take the log of a negative value. So this solution is `extraneous`, just ignore it. Only positive root works.

OpenStudy (wilder.monday):

thanks for the help. mind helping with a few more?

zepdrix (zepdrix):

Eh I'll be around :) Open up a new thread, I'll stop by if I can find time. You can @zepdrix and I'll try to find it.

OpenStudy (wilder.monday):

thanks :)